OpenStudy (et365):

(1+cos(y))/(1+sec(y)) in simplest numerical form?

OpenStudy (jdoe0001):

\(\bf \cfrac{1+cos(y)}{1+sec(y)}\qquad {\color{brown}{ sec(\theta)=\cfrac{1}{cos(\theta)}}}\quad thus\implies \cfrac{1+cos(y)}{1+\frac{1}{cos(y)}}\) what would that simplify to?

OpenStudy (et365):

That is where I hit the wall xD wondering what exactly to do with the denominator

OpenStudy (jdoe0001):

well \(\bf 1+\cfrac{1}{cos(y)}\implies \cfrac{1}{1}+\cfrac{1}{cos(y)}\implies \cfrac{}{{\color{brown}{ lcd?}}}\)

OpenStudy (jdoe0001):

what do you think the LCD or GCF would be?

OpenStudy (et365):

so cosy

OpenStudy (et365):

Now theres 1/cosy + 1/cosy?

OpenStudy (jdoe0001):

yeap... thus one sec

OpenStudy (jdoe0001):

\(\bf \cfrac{1+cos(y)}{1+sec(y)}\qquad {\color{brown}{ sec(\theta)=\cfrac{1}{cos(\theta)}}}\quad thus\implies \\ \quad \\ \cfrac{1+cos(y)}{\frac{cos(y)+1}{cos(y)}}\implies 1+cos(y)\cdot \cfrac{cos(y)}{cos(y)+1}\implies \cfrac{1+cos(y)}{1}\cdot \cfrac{cos(y)}{cos(y)+1} \\ \quad \\ \cfrac{\cancel{ cos(y)+1 }}{1}\cdot \cfrac{cos(y)}{\cancel{ cos(y)+1 }}\)

OpenStudy (jdoe0001):

hmm m issing a few ... anyhow \(\bf \cfrac{1+cos(y)}{1+sec(y)}\qquad {\color{brown}{ sec(\theta)=\cfrac{1}{cos(\theta)}}}\quad thus\implies \cfrac{1+cos(y)}{1+\frac{1}{cos(y)}} \\ \quad \\ \cfrac{1+cos(y)}{\frac{cos(y)+1}{cos(y)}}\implies 1+cos(y)\cdot \cfrac{cos(y)}{cos(y)+1}\implies \cfrac{1+cos(y)}{1}\cdot \cfrac{cos(y)}{cos(y)+1} \\ \quad \\ \cfrac{\cancel{ cos(y)+1 }}{1}\cdot \cfrac{cos(y)}{\cancel{ cos(y)+1 }}\)

OpenStudy (et365):

Thanks I see the step I was missing lol stay tuned I believe I have another question...should I make a new post or post it here?

OpenStudy (jdoe0001):

yes... thus if I dunno.. someone else may, and we can revise each other :)

OpenStudy (et365):

Gotcha xD soooo this one is basically the same thing but with (2+tan^2x)/(sec^2x) =(f(x))^2

OpenStudy (jdoe0001):

hmmm are you supposed to simplify \(\bf \cfrac{2+tan^2(x)}{sec^2(x)}?\)

OpenStudy (et365):

Yeah and that equals (f(x))^2

OpenStudy (jdoe0001):

yeap pretty much \(\bf \cfrac{2+tan^2(x)}{sec^2(x)}\implies \cfrac{2+\frac{sin^2(x)}{cos^2(x)}}{\frac{1}{cos^2(x)}}\implies \cfrac{\frac{2cos^2(x)+sin^2(x)}{cos^2(x)}}{\frac{1}{cos^2(x)}} \\ \quad \\ \cfrac{2cos^2(x)+sin^2(x)}{\cancel{ cos^2(x) }}\cdot \cfrac{\cancel{ cos^2(x) }}{1}\)

OpenStudy (et365):

LCD for the 2 in the numerator? I'm beginning to see thats where I stop thinking xD

OpenStudy (jdoe0001):

well the LCD atop is \(cos^2(x)\) so 2 is really 2/1 so \(cos^2(x) \div 1 = cos^2(x)\) then that * the 2 atop is \(2cos^2(x)\)

OpenStudy (et365):

Ahh, but now this answer needs some sort of numerical value or further simplification, I havent tried it with the f(x)^2

OpenStudy (et365):

Still no luck with that either, your results?

OpenStudy (jdoe0001):

hmm ok so \(\bf \cfrac{2+tan^2(x)}{sec^2(x)}\implies \cfrac{2+\frac{sin^2(x)}{cos^2(x)}}{\frac{1}{cos^2(x)}}\implies \cfrac{\frac{2cos^2(x)+sin^2(x)}{cos^2(x)}}{\frac{1}{cos^2(x)}} \\ \quad \\ \cfrac{2cos^2(x)+sin^2(x)}{\cancel{ cos^2(x) }}\cdot \cfrac{\cancel{ cos^2(x) }}{1}\implies 2cos^2(x)+sin^2(x) \\ \quad \\ cos^2(x)+cos^2(x)+sin^2(x)\qquad {\color{brown}{ sin^2(\theta)+cos^2(\theta)=1}}\qquad thus \\ \quad \\ cos^2(x)+1\)

OpenStudy (jdoe0001):

it could simplify to that

OpenStudy (et365):

Nope :\

OpenStudy (et365):

Not sure how to do this one lol

OpenStudy (jdoe0001):

hmm can you post a screenshot of the material?.... maybe it has some hint

OpenStudy (et365):

Yeah one sec

OpenStudy (jdoe0001):

hmmm the most I can think of it could come up to .... would be \(\bf cos^2(x)+1=[f(x)]^2\implies \pm \sqrt{cos^2+1}=f(x)\)

OpenStudy (et365):

Negative on both according to WeBWork lol

OpenStudy (jdoe0001):

ohhh ahemm smokes... just notice.... there's a -1 there

OpenStudy (et365):

You're right xD

OpenStudy (et365):

Its just cosx

OpenStudy (jdoe0001):

\( \bf \cfrac{2+tan^2(x)}{sec^2(x)}-1\implies \cfrac{2+\frac{sin^2(x)}{cos^2(x)}}{\frac{1}{cos^2(x)}}-1\implies \cfrac{\frac{2cos^2(x)+sin^2(x)}{cos^2(x)}}{\frac{1}{cos^2(x)}}-1 \\ \quad \\ \cfrac{2cos^2(x)+sin^2(x)}{\cancel{ cos^2(x) }}\cdot \cfrac{\cancel{ cos^2(x) }}{1}-1\implies 2cos^2(x)+sin^2(x)-1 \\ \quad \\ cos^2(x)+cos^2(x)+sin^2(x)\qquad {\color{brown}{ sin^2(\theta)+cos^2(\theta)=1}}\qquad thus \\ \quad \\ cos^2(x)\cancel{ +1-1 }\)

OpenStudy (jdoe0001):

well \(\bf cos^2(x)+cos^2(x)+sin^2(x)-1\qquad {\color{brown}{ sin^2(\theta)+cos^2(\theta)=1}}\qquad thus \\ \quad \\ cos^2(x)\cancel{ +1-1 }\)

OpenStudy (et365):

Just cosx the (f(x)) is squared so they were canceled

OpenStudy (jdoe0001):

and yes... once you take the square root... leaves you with cos(x) only

OpenStudy (et365):

This one seems easy its on the interval [0,2pi], I tried sqrt3/2 but no luck

OpenStudy (et365):

Ahh wait it has to be a multiple of pi

OpenStudy (jdoe0001):

\(\bf sin^2(t)=\cfrac{3}{4}\implies sin(t)=\sqrt{\cfrac{3}{4}}\implies sin^{-1}[sin(t)]=sin^{-1}\left(\cfrac{\sqrt{3}}{2}\right) \\ \quad \\ \measuredangle t=sin^{-1}\left(\cfrac{\sqrt{3}}{2}\right)\)

OpenStudy (et365):

which should be pi/3 correct?

OpenStudy (jdoe0001):

yes

OpenStudy (jdoe0001):

or 2pi/3

OpenStudy (et365):

Hmm let me include the whole problem and maybe youll get further than I

OpenStudy (jdoe0001):

k

OpenStudy (et365):

I've tried pi/3, 2pi/3, and 1/3

OpenStudy (et365):

Not sure how long you will be on, I'll be back in 5-15. Let me know what you get.

OpenStudy (jdoe0001):

your answer is correct, \(\frac{\pi}{3},\frac{2\pi}{3}\) are both correct HOWEVER the instructions ask to only give the rational, leaving the \(\pi\) out, since is already understood thus \(\bf \frac{1}{3},\frac{2}{3}\)

OpenStudy (et365):

I'm still here for now and yeah just tried lol nothing for that either >.>

OpenStudy (jdoe0001):

hmmm

OpenStudy (jdoe0001):

\(\bf sin^2(t)=\cfrac{3}{4}\implies sin(t)=\pm\sqrt{\cfrac{3}{4}}\implies sin^{-1}[sin(t)]=sin^{-1}\left(\cfrac{\pm\sqrt{3}}{2}\right) \\ \quad \\ \measuredangle t=sin^{-1}\left(\cfrac{\pm\sqrt{3}}{2}\right)\to \large \begin{cases} \frac{\pi}{3}\\ \frac{2\pi}{3}\\ \frac{4\pi}{3}\\ \frac{5\pi}{3} \end{cases}\)

OpenStudy (jdoe0001):

the \(\Large \pm\) of the root...would end up including all angles in each quadrant

OpenStudy (et365):

Ahh do + and - both?

OpenStudy (et365):

So*

OpenStudy (jdoe0001):

so, you'd need to include all those, the + ones and the - ones

OpenStudy (et365):

Yeah no luck lol not sure why but none